# Sagittarius A* near-infrared flares polarization as a probe of space-time I:

### v1

@@ 2.3 Gravastar
Similarly to relativistic Fluid spheres, gravitational vacuum stars, or Gravastars, are solutions of the Einstein’s field equations in the presence of an isotropic perfect fluid. In the case of Gravastars, this fluid is exotic, satisfying an equation of state of the form $p=-\rho$, where $p$ is the isotropic pressure and $\rho$ is the energy density. The interior of Gravastars is described by the metric
@@
where $\alpha$ is a parameter controlling the volumetric mass distribution of the star, and $m(r)=\frac{4}{3}\pi\rho r^{3}$ is the mass function. The exterior of the Gravastar is described by the Schwarzschild space-time. The parameter $\alpha$ is bounded between $1$ and a minimum value $\alpha_{\text{min}}$ that depends on the model, with $\alpha=1$ and $\alpha=\alpha_{\text{min}}$ corresponding to solutions with all mass distributed in volume or on the surface, respectively. Regardless of the mass distribution, these solutions develop a pair of bound-photon orbits whenever the radius $R$ of the surface of the Gravastar, where the matching between the interior and exterior space-times is performed, satisfies the condition $R\leq 3M$, with the limiting case $R=3M$ corresponding to a single degenerate bound-photon orbit. More details about this model can be found in Ref. Rosa et al. (2024). For $\alpha\neq 1$, the $g_{rr}$ coefficient of the metric is discontinuous, making the parallel transport of the polarization basis inside the ray tracing code GYOTO (more details in Section 3.1) impossible. We thus fixed $\alpha=1$ and selected three configurations: a model with $R=3M$ (GS1), $R=2.5M$ (GS2) and $R=2.01M$ (GS3). Again, we present the time integrated images of a hotspot orbiting these three configurations in Fig. 3.
Figure 3: Same as Fig. 2 with the Gravastar models with $R=2.01M$ in the top row, $R=2.5M$ in the middle row and $r=3M$ in the bottom row. Extracted from (Tamm et al., in prep).
3 Hot spot model
We consider an analytical orbiting hot spot model for the flares of Sgr A* in NIR as it is the model that agrees the most with the current data. However, we did not choose any specific physical phenomenon at the origin of the flare to be not model dependent. The hot spot is assumed to be a uniform sphere of plasma with a radius of $1\,r_{g}$ that emits synchrotron radiation.
@@ 3 Hot spot model
Figure 4: Astrometry of an orbiting hot spot in different metric models and simulated astrometric data with current GRAVITY uncertainties.
@@ 3.1 Polarized ray-tracing
We used the public polarized ray-tracing code GYOTO [^1] https://github.com/gyoto/Gyoto (Vincent et al., 2011; Aimar et al., 2024) to compute the images of the hot spot model for the four Stokes parameters (I, Q, U and V) which characterize the total received intensity (Stokes I) and the polarization of the received light (Q and U for linear polarization and V for circular polarization). For this study, we ignored the circular polarization V as it is not measurable by GRAVITY.
For each observing time, we compute the field integrated polarized fluxes $F_{I}(t)$, $F_{Q}(t)$, and $F_{U}(t)$ in Jansky, and the centroid position of the total flux ($X(t),Y(t)$) from the computed images. While polarization strongly depends on the orbital parameters and the inclination (see below), the incorporation of the astrometry allows breaking some degeneracies. For the rest of the paper, the field of view is set to 2.6 times the orbital radius in M units to optimize the computation time, and the default resolution is 300x300 pixels. Such resolution is sufficient to get the low-order plunge-through images from the ECO models but not the high-order ones, which require an extreme resolution but have a minor impact on the observed fluxes. The differences with the Schwarzschild metric come mainly from the additional or absence of light rings and additional plunge-through images (Rosa et al., 2025, Tamm et al., in prep). The observed wavelength is set to the one of GRAVITY, that is, $2.2\,\mu$m.
@@ 3.2 Orbital motion
For this theoretical study, we choose the simplest type of orbital motion, i.e., a circular orbit at radius $r$ in the equatorial plane with the Schwarzschild keplerian velocity, i.e., in Boyer-Lindquist coordinates,
$\theta$ $=\frac{\pi}{2}$ (4)
$\frac{dr}{dt}$ $=\frac{d\theta}{dt}=0$ (5)
$\frac{d\varphi}{dt}$ $=r^{-3/2}.$ (6)
Note that in the ECO metrics, this choice of velocity does not necessarily correspond to time-like geodesics, nor to the Keplerian velocity in these metrics. However, this study is focused on the detectability of the effect of alternative metrics on the polarization signals. Thus, to make a proper comparison, the 3D-velocity must be the same in all metrics. Nevertheless, this assumption does not change our conclusions, as the change in velocity is negligible for non-rotating ECOs compared to the typical error bars of the problem. Furthermore, for orbital radius larger than the size of the Boson star, which is the case here, the difference of velocity between the Boson star metric and Schwarzschild is very small and can be neglected. For Gravastars and Fluid spheres, the velocity is identical to Schwarzschild outside the compact object radius.
@@ 3.3 Electron Energy Distribution function
The IR flux from Sgr A* flares is thought to be generated by synchrotron radiation from a non-thermal population of electrons (GRAVITY Collaboration et al., 2021). We thus consider a $\kappa$-distribution of electrons (Marszewski et al., 2021) for our hot spot. This distribution is characterized by a thermal core with a temperature $T_{e}$ and a power-law tail with $\kappa$-index222related to the power law index $p$ as $\kappa=p+1$ and a number density $n_{e}$. These three parameters have a strong influence on the fluxes, while the $\kappa$-index governs the NIR spectral index. To simplify and limit the number of parameters, we set the number density to $5\times 10^{6}$ cm-3, the temperature to $5\times 10^{10}$ and the $\kappa$-index to $3.5$. The overall flux of the generated synthetic flares with these settings matches the usual observed flux from Sgr A* flares ($\approx 10$ mJy). Obviously, not all flares have the same maximum flux. To mitigate this, we restricted the study to the normalized polarized fluxes $Q/I$ and $U/I$.
@@ 3.4 Magnetic field configuration
The average of polarization measurements of multiple flares made in Gravity Collaboration et al. (2023) has shown that the magnetic field configuration in the flare emitting region is vertical. Indeed, the polarization angular velocity of $\sim 6$°/min is only compatible with the vertical configuration of the magnetic field lines at low inclination. This statement remains true even in the ECOs metrics (Rosa et al., 2025, Tamm et al., in prep). We thus restrict our study to this magnetic field configuration. The magnetic field strength is defined through the magnetization parameter $\sigma=B^{2}/4\pi m_{p}c^{2}n_{e}$ fixed to 0.01. The magnetic field strength is thus around 30 G which is the expected order of magnitude for Sgr A* flares (von Fellenberg et al., 2025). Note that in this model, this parameter is only a scaling factor of the total flux.
The observed linear polarization fraction of Sgr A* flares is $\sim 20-25\%$, much lower than the linear polarization fraction expected from a purely ordered magnetic field. A probable explanation is that the magnetic field is only partially ordered (Event Horizon Telescope Collaboration et al., 2024). The contribution of plunge-through images can also result in a decrease of the linear polarization fraction, but only for a part of the orbit and cannot be responsible for the global difference of polarization fraction. To take into account that the magnetic field is only partially ordered, we apply a constant scaling factor $lp$ to the polarized quantities (Stokes Q and U), whose value is in [0,1] [^3] This parameter should not be interpreted as the degree of order of the magnetic field as the relation is indirect and non-trivial..
Our model thus has a total of five parameters, summarized in Tab. 1, four for the hot spot: orbital inclination $i$, radius $r$, initial azimuthal angle $\varphi_{0}$ and the polarization factor $lp$ and one for the metric (metric model). We fixed the Position Angle of the Line of Nodes to 177.3 following Gravity Collaboration et al. (2023) as this parameter add an unnecessary degree of freedom for this theoretical study.
@@
Linear polarization factor $lp$ $0.4$
@@ 4 Methodology
To determine the detectability of the metrics considered in this paper, we performed fits of simulated data with all the metric and compared the results of these fits. Using statistical criteria, we conclude on the detectability of the metrics. The following section describes in detail our methodology.
@@ 4.1 Generation of simulated data
We generated simulated data following the set of parameters presented in Tab. 1 for the hot spot in each ECO model metric.
In order to accommodate the uncertainties inherent in observational data, we introduce a random variable to the simulated data points, which is drawn from a Gaussian distribution. For the simulated data on astrometric measurements, the Gaussian distribution is characterized by a standard deviation denoted $\sigma_{\mathrm{Astro}}=30\,\mu$as. Similarly, for the simulated polarized flux measurements (I, Q and U), the standard deviation is represented by $\sigma_{\mathrm{Flux}}$. Sect. 5 provides different uncertainty values $\sigma_{i}$, aligning with GRAVITY’s actual uncertainties ($\sigma_{\mathrm{Flux}}=0.2$ mJy) and the improved precisions anticipated for GRAVITY+ upgrade ($\sigma_{\mathrm{Flux}}=0.03$ mJy).
The observational duration and data point intervals are by default calibrated to resemble those of GRAVITY, namely, one measurement every 5 minutes resulting in a total of 12 observations (amounting to 60 minutes overall). In Sect. 5, we also investigated the impact of a better time sampling on the detectability of the metrics.
@@ 4.2.1 Fitting procedure
Contrary to the four hot spot parameters, the metric parameter is not continuous (a range of values) but is instead a choice of configuration. The treatment of this parameter is thus different from the others. As part of our fit strategy, we opted to individually fit the four remaining parameters of our model, namely, the inclination, the orbital radius, the initial azimuthal angle and the polarization factor, within the context of all possible ECO + Schwarzschild background metrics. Subsequently, we conducted a comparative analysis of the results obtained.
We perform the fits in two distinct steps: 1) a grid, 2) a follow-up of the $\chi^{2}$ gradient with fixed steps. For each background metric, we computed a grid of modeled data varying the orbital radius, the inclination, and the initial azimuthal angle. The linear polarization fraction is a simple scaling factor and can thus be applied as post-processing and does not require the use of GYOTO. The minimum and maximum values as the number of grid points of each parameter (including the linear polarization factor) are listed in Tab. 2. For each point on the grid, the modeled observable quantities ($Q/I(t)$, $U/I(t)$, $X(t)$, $Y(t)$) are estimated from the simulated model using linear interpolation. We evaluate the $\chi^{2}_{\mathrm{red}}$ as
$\chi^{2}_{\mathrm{red}}=\frac{1}{\mathrm{\mathrm{dof}}}\sum_{i}\left(\frac{x_{i}-\mu_{i}}{\sigma_{i}}\right)^{2},$ (7)

### v3

@@ 2.3 Gravastar
Similarly to relativistic Fluid spheres, gravitational vacuum stars, or Gravastars, are solutions of the Einstein’s field equations in the presence of an isotropic perfect fluid. In the case of Gravastars, this fluid is exotic, satisfying an equation of state of the form $p=-\rho$, where $p$ is the isotropic pressure and $\rho$ is the energy density (Mazur and Mottola, 2004; Visser and Wiltshire, 2004; Mottola and Vaulin, 2006; Pani et al., 2009; Mazur and Mottola, 2015; Danielsson and Giri, 2018; Posada and Chirenti, 2019; Mazur and Mottola, 2023). The interior of Gravastars is described by the metric
@@
where $\alpha$ is a parameter controlling the volumetric mass distribution of the star, and $m(r)=\frac{4}{3}\pi\rho r^{3}$ is the mass function. The exterior of the Gravastar is described by the Schwarzschild space-time. The parameter $\alpha$ is bounded between $1$ and a minimum value $\alpha_{\text{min}}$ that depends on the model, with $\alpha=1$ and $\alpha=\alpha_{\text{min}}$ corresponding to solutions with all mass distributed in volume or on the surface, respectively. Regardless of the mass distribution, these solutions develop a pair of bound-photon orbits whenever the radius $R$ of the surface of the Gravastar, where the matching between the interior and exterior space-times is performed, satisfies the condition $R\leq 3M$, with the limiting case $R=3M$ corresponding to a single degenerate bound-photon orbit. More details about this model can be found in Ref. Rosa et al. (2024). For $\alpha\neq 1$, the $g_{rr}$ coefficient of the metric is discontinuous, making the parallel transport of the polarization basis inside the ray tracing code GYOTO (more details in Section 3.1) impossible. We thus fixed $\alpha=1$ and selected three configurations: a model with $R=3M$ (Gravastar 1), $R=2.5M$ (Gravastar 2) and $R=2.01M$ (Gravastar 3). Again, we present the time integrated images of a hot-spot orbiting these three configurations at low inclination in Fig. 2. All configurations present a set of light-rings and at least one plunge through image. The size of the inner plunge-through image decreases with an increase in compacticity, with the most compact configuration (Gravastar 3) presenting an additional light ring.
Figure 2: Time integrated image of a hot-spot orbiting the Gravastar models with $R=3M$ (Gravastar 1) in the left panel, $R=2.5M$ (Gravastar 2) in the middle panel and $r=2.01M$ (Gravastar 3) in the right panel. Extracted from Tamm et al. (2025).
Metric model R $r_{g}$] Compactness # of light ring(s)
Schwarzschild 2 0.5 1
Boson star 2 3.95 $\sim 0.253$ 1
Boson star 3 3.12 $\sim 0.321$ 2
Fluid sphere 2 2.5 0.4 2
Fluid sphere 3 3 $\sim 0.333$ 2 (degen.)
Gravastar 1 3 $\sim 0.333$ 2 (degen.)
Gravastar 2 2.5 0.4 2
Gravastar 3 2.01 $\sim 0.498$ 2
Table 1: Summary of the metric models properties.
3 Hot-spot model
We consider an analytical orbiting hot-spot model for the flares of Sgr A* in NIR as it is the model that agrees the most with the current data. However, we did not choose any specific physical phenomenon at the origin of the flare to avoid model dependence. The hot-spot is assumed to be a uniform sphere of plasma with a radius of $1\,r_{g}$ that emits synchrotron radiation.
@@ 3 Hot-spot model
Figure 3: Astrometry of an orbiting hot-spot in different metric models and simulated astrometric data with current GRAVITY uncertainties.
@@ 3.1 Polarized ray-tracing
We used the public polarized ray-tracing code GYOTO [^2] https://github.com/gyoto/Gyoto (Vincent et al., 2011; Aimar et al., 2024) to compute the images of the hot-spot model for the four Stokes parameters (I, Q, U and V) which characterize the total received intensity (Stokes I) and the polarization of the received light (Q and U for linear polarization and V for circular polarization). For this study, we ignored the circular polarization V as it is not measurable by GRAVITY.
[…]
@@ 3.2 Orbital motion
For this theoretical study, we choose the simplest type of orbital motion, i.e., a circular orbit at radius $r$ in the equatorial plane. The velocity of Sgr A*’s IR flares as well as the spin of Sgr A* are not well constrained because of the large astrometric uncertainties. Some recent studies are suggesting a possible super-Keplerian motion (GRAVITY Collaboration et al., 2018b; Aimar et al., 2023; Antonopoulou and Nathanail, 2024; Yfantis et al., 2024a, b; Xie et al., 2025). Moreover, as shown in Fig. 3, the astrometric tracks in the ECOs metric are smaller for the same orbital radius due to the contribution of the inner plunge-through images. Thus, when we want to fit simulated data of a Keplerian orbit in an ECO metric (see more details in Sect. 4) using the Kerr metric, the apparent orbital radius will be smaller, and the velocity will be sub-Keplerian. On the contrary, when fitting simulated data of a keplerian orbit in Kerr using an ECO metric, the apparent orbit will be larger with a super-Keplerian motion. Thus, we allow the orbital velocity to be a free parameter in the form of a coefficient $K_{\mathrm{coef}}$ which is the fraction of the Schwarzschild Keplerian velocity. The orbital motion is thus defined in Boyer-Lindquist coordinates as
$\theta$ $=\frac{\pi}{2},$ (4a)
$\frac{dr}{dt}$ $=0,$ (4b)
$\frac{d\theta}{dt}$ $=0,$ (4c)
$\frac{d\varphi}{dt}$ $=K_{\mathrm{coef}}\cdot\left(r^{3/2}+a\right)^{-1},$ (4d)
with $a$ the spin of the compact object (only for the Kerr metric in this study).
@@ 3.3 Electron Energy Distribution function
The IR flux from Sgr A* flares is thought to be generated by synchrotron radiation from a non-thermal population of electrons (GRAVITY Collaboration et al., 2021). We thus consider a $\kappa$-distribution of electrons (Marszewski et al., 2021) for our hot-spot. This distribution is characterized by a thermal core with a temperature $T_{e}$ and a power-law tail with $\kappa$-index333related to the power law index $p$ as $\kappa=p+1$ and a number density $n_{e}$. These three parameters have a strong influence on the fluxes, while the $\kappa$-index governs the NIR spectral index. To simplify and limit the number of parameters, we set the number density to $10^{7}$ cm-3, the temperature to $5\times 10^{10}$K and the $\kappa$-index to $3.5$. The overall flux of the generated synthetic flares with these settings matches the usual observed flux from Sgr A* flares ($\approx 10$ mJy). Obviously, not all flares have the same maximum flux. To mitigate this, we restricted the study to the normalized polarized fluxes $Q/I$ and $U/I$.
@@ 3.4 Magnetic field configuration
The average of polarization measurements of multiple flares made by GRAVITY shows a single loop of polarization over one orbital period of P=60P=60\,min which constrains both the magnetic field configuration to be vertical and the inclination to $i=157\degree\pm 5\degree$ (Gravity Collaboration et al., 2023, see Fig. 4 of). This magnetic field configuration has also been investigated for ECO-metrics (Rosa et al., 2025; Tamm et al., 2025). We thus restrict our study to this magnetic field configuration. The magnetic field strength is defined through the magnetization parameter $\sigma=B^{2}/4\pi m_{p}c^{2}n_{e}$ fixed to 0.01. The magnetic field strength is thus around 43 G which is the expected order of magnitude for Sgr A* flares (von Fellenberg et al., 2025). Note that in this model, this parameter is only a scaling factor of the total flux.
The observed linear polarization fraction of Sgr A* flares is $\sim 10-40\%$ (GRAVITY Collaboration et al., 2020c), much lower than the linear polarization fraction expected from a purely ordered magnetic field (without any stochastic component). A probable explanation is that the magnetic field is only partially ordered, i.e. with a stochastic component, as suggested by the EHT observation of Sgr A* (Event Horizon Telescope Collaboration et al., 2024). This effectively reduces the observed polarization fraction (Rybicki and Lightman, 1979). The contribution of plunge-through images can also result in a decrease of the linear polarization fraction, but only for a specific part of the orbit and cannot be responsible for the global difference of polarization fraction. To take into account the low observed polarization fraction compared to the models, we apply a constant scaling factor $lp$ to the modeled polarized quantities (Stokes Q and U) from GYOTO, whose value is in [0,1] [^4] This parameter should not be interpreted as the degree of order of the magnetic field as the relation is indirect and non-trivial..
3.5 Parameters summary
Our model has a total of six free parameters. The hot-spot has five free parameters, summarized in Tab. 2: inclination $i$, orbital radius $r$, initial azimuthal angle $\varphi_{0}$, orbital velocity through the fraction of the Keplerian velocity coefficient $K_{\mathrm{coef}}$ and the polarization factor $lp$. The last parameter is the background metric (metric model). We fixed the Position Angle of the Line of Nodes to 177.3 following Gravity Collaboration et al. (2023) as this parameter add an unnecessary degree of freedom for this theoretical study.
The effect of the spin of a Kerr black hole on the polarization measurements $Q/I$ and $U/I$ could be similar to the impact of the plunge-through images from the ECOs metrics. Thus, to test if an ECO signature can be fitted and thus mimic by the spin of a Kerr black hole, we add the spin as a free parameter in the Kerr metric (see 5.2).
@@ 3.5 Parameters summary
Orbital velocity scaling factor $K_{\mathrm{coef}}$ 1.0
Linear polarization factor lplp $0.4$
@@ 4 Methodology
To determine the detectability of the metrics considered in this paper, we performed fits of simulated data with all aforementioned metrics and compared the results of these fits. Using statistical criteria, we conclude on the detectability of the metrics. The following section describes in detail our methodology.
@@ 4.1 Generation of simulated data
We generated simulated data following the set of parameters presented in Tab. 2 for the hot-spot within a given background metric.
[…]
$\sigma_{\mathrm{Flux}}^{\mathrm{GRAVITY}}=0.15\cdot I^{0.6},$ (5)
with $\sigma_{\mathrm{Flux}}^{\mathrm{GRAVITY}}$ and $I$ in mJy. The flux dependency of Eq. 5 has been reported by GRAVITY Collaboration et al. (2020a) and here calibrated with the uncertainties from (Gravity Collaboration et al., 2023) [^5] The uncertainties in (Gravity Collaboration et al., 2023) are two times smaller than the one predicted by GRAVITY Collaboration et al. (2020a). We thus applied a factor 0.5 to their fitted equation.. The uncertainties associated with the ratios $Q/I$ and $U/I$ are derived by propagating the measurement errors in $I$, $Q$, and $U$. In Sect. 5 we investigate the detectability of the ECOs with various flux uncertainties in the form of fraction of the Eq. 5, from 1 being the GRAVITY’s actual uncertainties until $\sigma_{\mathrm{Flux}}^{\mathrm{GRAVITY}}/7$ which correspond to the best anticipated improved precisions for GRAVITY+ upgrade (Bourdarot and Eisenhauer, 2024).
The observational duration and data point intervals are by default calibrated to resemble those of GRAVITY, namely, one measurement every 5 minutes resulting in a total of 12 observations, amounting to 60 minutes overall equivalent to one full orbital period. In Sect. 5, we also investigated the impact of a better time sampling on the detectability of the metrics.
@@ 4.2.1 Fitting procedure
Contrary to the five hot-spot parameters, the metric parameter is not continuous (a range of values) but is instead categorical, i.e., a choice of a configuration. The treatment of this parameter is thus different from the others. As part of our fit strategy, we opted to individually fit the five (six for Kerr) remaining parameters of our model, namely, the inclination, the orbital radius, the initial azimuthal angle, the orbital velocity scaling factor and the polarization factor (plus the spin in Kerr), within the context of all possible ECO + Kerr background metrics.
We perform a classical $\chi^{2}$ minimization using the scipy.least_square algorithm with the Trust Region Reflective method. This combination is selected because it is well suited to large, bounded optimization problems, which corresponds to our situation involving five or even six parameters with clearly defined limits. Another important aspect of the fitting strategy is the computational cost. Each model evaluation requires several minutes, so a single fit typically ranges from about 2 to 12 hours, depending on the chosen metric. This type of algorithm is highly dependent on the initial guess. To reduce the risk of converging to a local minimum, we adopt as our first-guess the point with the minimum reduced chi-square, $\chi^{2}_{\mathrm{red}}$, obtained from a preliminary grid search (see Appendix A for details on the grid). We evaluate the $\chi^{2}_{\mathrm{red}}$ as
$\chi^{2}_{\mathrm{red}}=\frac{1}{\mathrm{\mathrm{\nu}}}\sum_{i}\left(\frac{x_{i}-\mu_{i}}{\sigma_{i}}\right)^{2},$ (6)